OpenStudy (lakers):

Classify the triangle based on the side lengths 9, 40 and 41.

OpenStudy (anonymous):

The triangle with no equal sides is called Scalene Triangle.. The triangle two sides equal is called Isosceles Triangle.. The triangle with all the three sides equal is called Equilateral Triangle..

OpenStudy (lakers):

is it a right, acute, obtuse, or no triangle?

OpenStudy (zeesbrat3):

without the angle measurements it is impossible to tell

OpenStudy (lakers):

Classify the triangle based on the side lengths 9, 40 and 41. A. right B. acute C. obtuse D. no triangle can be formed with given side lengths

OpenStudy (lakers):

this is the whole question

OpenStudy (zeesbrat3):

If it is a right triangle then the lengths will equal out in pythagorean theorem

OpenStudy (lakers):

so is it no triangle?

OpenStudy (zeesbrat3):

it is a triangle

OpenStudy (zeesbrat3):

I would say right because using pythagorean theorem, it comes out equal

OpenStudy (zeesbrat3):

\[9^2+40^2=41^2 81+1600=1681\]

OpenStudy (anonymous):

Oh I did not check that.. \[\sqrt{1600 + 81} = 41\]

OpenStudy (anonymous):

Yes as pythagoras theorem is applicable here so It is A Right Triangle..

OpenStudy (lakers):

ok thanks guys i posted another 1 like this if yu can help me

OpenStudy (anonymous):

Welcome dear..